OpenStudy (viixxen):

Simplify the rational expression by rationalizing the denominator. 4√6/√30 Thank you! :)

OpenStudy (muzzack):

well 30 you have to do the prime factorization

OpenStudy (muzzack):

it would be √5*2*3

OpenStudy (johnweldon1993):

\[\large \frac{4\sqrt{6}}{\sqrt{30}} \times \frac{\sqrt{30}}{\sqrt{30}} = \frac{4\sqrt{6}\sqrt{30}}{30} = \frac{4\sqrt{180}}{30} = \frac{4\sqrt{36 \times 5}}{30} = \frac{24\sqrt{5}}{30} = \frac{4\sqrt{5}}{5}\]